Lauryl alcohol is obtained from coconut oil and is used to make detergents. A solution of 5.00 g of lauryl alcohol in 0.100 kg of benzene freezes at 4.1oC. What is the molar mass of this substance?
1
Expert's answer
2020-12-14T14:46:22-0500

Raoult's law states:

"T_b - T_s = KC_m", where Tb -- freezing point of benzene (5.5oC), Ts -- freezing point of the solution, K -- cryoscopic constant (5.12 K*kg/mol for benzene), Cm -- molality of solution.

"C_m = m_l\/Mm_s" , where ml -- mass of lauryl alcohol, ms -- mass of the solvent.

Combine the two formulas:

"T_b - T_s = Km_l\/Mm_s"

Find the molar mass:

"M = Km_l\/m_s(T_b-T_s)"

"M = 5.12*5\/0.1(5.5-4.1) =" 183 (g/mol)

Answer: 183 g/mol -- the molar mass of lauryl alcohol. (The real one is 186 g/mol).